Determine the polynomial function whose graph passes through the points (2, 4), (3,6) and (5,10). Also sketch the graph of the polynomial function. (Using Cramers Method).
1
Expert's answer
2020-12-15T07:03:27-0500

These tree points lie on the straight line "y=2x". The power of polynomial is 1. If we will find polynomial with power 2 as "y=ax^2+bx+c" we will

receive "a=0, b=2, c=0", because parabola can have only two roots,

and "y=2x":

"\begin{cases}
4a+2b+c=4\\
9a+3b+c=6\\
25a+5b+c=10
\end{cases}"


"a=det\begin{Vmatrix}
 4 & 2 & 1 \\
 6 & 3 & 1\\
 10&5 &1
\end{Vmatrix} / 
det\begin{Vmatrix}
 4 & 2 & 1\\
 9 & 3 & 1\\
 25 & 5 & 1
\end{Vmatrix}
=" "=(4(3-5)-2(6-10)+1(6*5-3*10))\/"

"(4(3-5)--2(9-25)+1(9*5-25*3))="

"=(-8+8+0)\/(-8+32-30)=0\/(-6)=0"


"b=det\begin{Vmatrix}
 4 & 4 & 1 \\
 9 & 6 & 1\\
 25&10 &1
\end{Vmatrix} / 
det\begin{Vmatrix}
 4 & 2 & 1\\
 9 & 3 & 1\\
 25 & 5 & 1
\end{Vmatrix}
="

"=(4(6-10)-4(9-25)+(9*10-25*6))\/(-6)="

"=(-16+64-60)\/(-6)=(-12)\/(-6)=2"


"c=det\begin{Vmatrix}
 4 & 2 & 4 \\
 9 & 3 & 6\\
 25&5 &10
\end{Vmatrix} / 
det\begin{Vmatrix}
 4 & 2 & 1\\
 9 & 3 & 1\\
 25 & 5 & 1
\end{Vmatrix}
="

"=(4(3*10-5*6)-2(9*10-6*25)+"

"+4(9*5-3*25))\/(-6)="

"=(4*0+2*60-4*30)\/(-6)=0"


If we want to find polynomial with power 3, we need solve system:

"\begin{cases}
8a+4b+2c+d=4\\
27a+9b+3c+d=6\\
125a+25b+5c+d=10
\end{cases}"

and this system will have infinitely many solutions.